# Rigidity of anti-de Sitter (2+1)-spacetimes with convex boundary near the Fuchsian locus

### v1

@@ 4.2. Local infinitesimal rigidity
Similarly we define a differential operator $\mathcal{L}$. Tamburelli showed that for all such $Z^{\prime}$ the operators $\mathcal{L}_{i}$, $\mathcal{L}$, as operators from $C^{\infty}(Z^{\prime})$ to $C^{\infty}(\Lambda^{2}TS\otimes TS)$, are elliptic of index $66k-$ 6 with trivial cokernel, where $k$ is the Euler characteristic of $S$. See the proof of Lemma 3.2 in [59]. (Compare also with [40, Lemma 3.1] of Labourie.) Hence every $\mathcal{L}_{i}$ on $C^{\infty}(Z_{i})$ has kernel of dimension $66k-$ 6, where $C^{\infty}(Z_{i})$ is the tangent space to $C^{\infty}(J)$ at $j_{i}$. We denote this kernel by $X_{i}$. The same holds for the kernel $X$ of $\mathcal{L}$ at $C^{\infty}(Z)$, the tangent space at $j$.
We claim that it is enough to show that $X_{i}$ converge to $X$ as subspaces of $\mathcal{A}$. Indeed, since $\mathcal{K}_{i}$ are isomorphisms converging to $\mathcal{K}$, if $X_{i}$ converge to $X$, then $T_{b_{i}}\mathcal{GC}(s_{i})=\mathcal{K}_{i}(X_{i})$ converge to $T_{b}\mathcal{GC}(s)=\mathcal{K}(X)$. To show the convergence of $X_{i}$, we construct a sequence of bundle automorphisms of $T^{1}_{1}S$ converging to the identity that send $Z_{i}$ to $Z$. To this purpose, we show that there exists a sequence of projectivized operator fields $a_{i}\in C^{\infty}(\mathbb{P}(T^{1}_{1}S))$, $C^{\infty}$-converging to the identity, such that $a_{i}j_{i}a_{i}^{-1}=j$. Here we note that the conjugation operation is well-defined for invertible elements of the projectivized operator bundle.
For the moment, let us consider a vector space $V\cong\mathbb{R}^{2}$ with some orientation. Let $J$ be the space of orientation-preserving almost-complex structures on $V$. Pick $j,j_{1}\in J$. We claim that there exists a canonical $a\in{\rm PSL}(2,\mathbb{R})$ such that $aj_{1}a^{-1}=j$, where by canonical we mean that it is independent on a choice of basis or on a choice of any additional structure. Indeed, under the identification ${\rm PSL}(2,\mathbb{R})\cong\mathbb{A}d\mathbb{S}^{3}$ the space $J$ becomes a connected component of the set of points at the timelike unit distance from the identity. Under the identification
@@
the diagonal is the set of isometries fixing the identity. Hence, there exists a unique isometry fixing the identity, sending $j_{1}$ to $j$ and preserving the geodesic between them. This isometry is given by conjugation by $a\in{\rm PSL}(2,\mathbb{R})$.
Now we return to our setting. For every $i$ at every point of $S$ we pick the respective element of the fiber of $\mathbb{P}(T^{1}_{1}S)$ conjugating $j_{i}$ to $j$. This produces a smooth section $a_{i}$ of the projectivized bundle $\mathbb{P}(T^{1}_{1}S)$ such that $a_{i}j_{i}a_{i}^{-1}=j$.
Denote now $L^{2}_{1}(Z)$ by $\mathcal{Z}$. The conjugation by the projectivized operator fields $a_{i}$ define isomorphisms $\mathcal{R}_{i}:\mathcal{A}\rightarrow\mathcal{A}$ that converge to the identity in the sense of operator norms. The operators $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ are Fredholm operators from $\mathcal{Z}$ to $L^{2}(\Lambda^{2}TS\otimes TS)$ that converge to a Fredholm operator $L$. All $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ and $\mathcal{L}$ have the same dimension of kernel and zero cokernel. Denote the kernels of $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ by $Y_{i}$. For every $a\in\mathcal{A}$ we have
$d(a,X)\gamma(\mathcal{L})\leq\|\mathcal{L}a\|.$
It follows that for $a\in Y_{i}$, $\|a\|\leq 1$, we get
$d(a,X)\leq\gamma^{-1}(\mathcal{L})\|(\mathcal{L}-(\mathcal{R}_{i})_{*}\mathcal{L}_{i})\|.$
@@
We can now prove Lemma 4.2.
Proof of Lemma 4.2..
@@
By a combination of Lemma 4.4 and Lemma 4.1, these sequences subconverge to $\dot{b}_{+}\in T_{b_{+}}\mathcal{GC}_{+}(s_{+})$ and $\dot{b}_{-}\in T_{b_{-}}\mathcal{GC}_{-}(s_{-})$. Due to the definition of the maps ϕs±,±subscriptitalic-$\phi_{s_{\pm},\pm}$ via formula (2.4), equation (4.2) implies that
@@ 4.3. Local homeomorphism
Now our goal is to prove the following statement.
Lemma 4.5.
Let $e\in\mathcal{E}_{F}$. There exists its neighborhood $U_{e}$ in $\mathcal{E}$ such that the image of $U_{e}$ by $\mathcal{I}$ is open and $\mathcal{I}|_{U_{e}}$ is a homeomorphism onto its image.
We need a preliminary work. Let $\Sigma$ be a spacelike surface in $\mathbb{A}d\mathbb{S}^{3}$, $n$ be the field of future-oriented unit normals, and $u$ be a vector field over $\Sigma$, which is decomposed as $u=v+an$ for a tangent vector field $v$ and a function $a$. Denote the first and the second fundamental forms of $\Sigma$ by $\mathrm{I}$ and $\mathrm{I\!I}$ respectively, and denote by $\nabla$ the induced connection on $\Sigma$. It is a standard computation, to check that the derivative of $\mathrm{I}$ along $u$ is given by (see, e.g., [30])
@@
The bundle $Z$ of symmetric $(0,2)$-tensors over $\Sigma$ has rank 3. Denote by $Z_{1}$ its subbundle of tensors orthogonal to $\mathrm{I\!I}$ with respect to $\mathrm{I\!I}$. Furthermore, let $b$ be the shape operator, $j$ be the almost-complex structure of $\mathrm{I\!I}$ and $\nabla^{\mathrm{I\!I\!I}}$ be the Levi–Civita connection of the third fundamental form. Define a differential operator $\mathcal{P}:C^{\infty}(T\Sigma)\rightarrow C^{\infty}(Z)$ by
@@
A direct computation (see, e.g., [54, Proposition 8.4]) shows that its image is in $C^{\infty}(Z_{1})$. Furthermore, it was proven in [54, Proposition 8.4] that
@@
When $f:\widetilde{S}\rightarrow\mathbb{A}d\mathbb{S}^{3}$ is an equivariant spacelike embedding, we can pullback $\mathcal{P}$ to $S$. It has the same principal symbol as a similarly defined operator, obtained by replacing $\nabla^{\mathrm{I\!I\!I}}$ with $\nabla^{\mathrm{I\!I}}$, the Levi–Civita connection of $\mathrm{I\!I}$. The latter operator has the same kernel and cokernel as the standard Cauchy–Riemann operator $\bar{\partial}$ (defined on de Rham complex tensored with $TS$). It follows from the Atiyah–Singer theorem (see, e.g., [32, Section 2]) that
@@
For a closed surface $S$ of genus $k$, as an operator $C^{\infty}(TS)\rightarrow C^{\infty}(Z_{1})$, where the latter is now considered as a bundle over $S$, $\mathcal{P}$ is an elliptic operator of index $-(6k-6)$.
Now we make another step aside. For the moment, let $e$ be an element of $\mathcal{E}_{+}$. Consider its lift $(\rho,f)\in\widehat{\mathcal{E}}_{+}$. Pick a right inverse $\mu:H^{1}_{\rho}(\mathfrak{g})\rightarrow Z^{1}_{\rho}(\mathfrak{g})$ to the projection $Z^{1}_{\rho}(\mathfrak{g})\rightarrow H^{1}_{\rho}(\mathfrak{g})$. Recall that $\widetilde{V}_{f}$ is the restriction of $T\mathbb{A}d\mathbb{S}^{3}$ to $f(\widetilde{S})$, which we now denote just by $\widetilde{V}$, and that $C^{\infty}_{\rho}(\widetilde{V})$ is the space of automorphic vector fields with respect to $\dot{\rho}\in Z^{1}_{\rho}(\mathfrak{g})$. Denote by $C^{\infty}_{\rho,n}(\widetilde{V})$ the subspace of automorphic vector fields with respect to $\dot{\rho}$ in the image of $\mu$. We can identify $T_{e}\mathcal{E}_{+}\cong C^{\infty}_{\rho,n}(\widetilde{V})$. Denote by $V$ the projection of $\widetilde{V}$ to $S$. There is a natural projection $C^{\infty}_{\rho,n}(\widetilde{V})\rightarrow H^{1}_{\rho}(\mathfrak{g})$, whose kernel consists of equivariant vector fields, and hence can be identified with $C^{\infty}(V)$. We pick a right inverse $\nu:H^{1}_{\rho}(\mathfrak{g})\rightarrow C^{\infty}_{\rho,n}(\widetilde{V})$. The choices of $\mu$ and $\nu$ allow to identify
(4.5) $T_{e}\mathcal{E}_{+}\cong H^{1}_{\rho}(\mathfrak{g})\times C^{\infty}(V).$
Consider now a neighborhood $U$ of $e$ and a chart $\phi:U\rightarrow T_{e}\mathcal{E}_{+}$ adapted for the projection $\pi_{+}:\mathcal{E}_{+}\rightarrow\mathcal{R}$. Here “adapted” means that for any $e^{\prime}$ and any $\tau\in H^{1}_{\rho}(\mathfrak{g})$ the composition
@@
is nonzero. It is possible to choose such a chart, since $\pi_{+}$ is a submersion. We get an identification
@@
Every $e^{\prime}\in U$ defines a decomposition of $V$ into two subbundles, the tangent one, and the normal one (with respect to the metric on $V$ induced by $e^{\prime}$). We denote them by $V_{e^{\prime},1}$ and $V_{e^{\prime},2}$. Also define $V_{1}:=V_{e,1}$ and $V_{2}:=V_{e,2}$. Let $\alpha_{e^{\prime},1}:V_{e^{\prime},1}\rightarrow V_{1}$ and $\alpha_{e^{\prime},2}:V_{e^{\prime},2}\rightarrow V_{2}$ be the orthogonal projections with respect to the metric on $V$ induced by $e$.
Also every $e^{\prime}\in U$ defines a decomposition of $Z$ into two subbundles, $Z_{e^{\prime},1}$ and $Z_{e^{\prime},2}$, where $Z_{e^{\prime},2}$ is spanned by $\mathrm{I\!I}_{e^{\prime}}$, the second fundamental form induced on $S$ by $e^{\prime}$, and $Z_{e^{\prime},1}$ is $\mathrm{I\!I}_{e^{\prime}}$-orthogonal to it. Define $Z_{1}:=Z_{e,1}$ and $Z_{2}:=Z_{e,2}$. Let $\beta_{e^{\prime},1}:Z_{e^{\prime},1}\rightarrow Z_{1}$ and $\beta_{e^{\prime},2}:Z_{e^{\prime},2}\rightarrow Z_{2}$ be the orthogonal projections with respect to the metric on $Z$ induced by $\mathrm{I\!I}_{e}$. For the map $\xi:U\times H^{1}_{\rho}(\mathfrak{g})\rightarrow C^{\infty}(Z)$ defined above, denote its projections to $Z_{1}$ and to $Z_{2}$ by $\xi_{1}$ and $\xi_{2}$.
Now, however, we have to slightly change all this notation. Namely, we now pick $e\in\mathcal{E}$ rather than from $\mathcal{E}_{+}$. Then for some its neighborhood $U$ decomposition (4.6) will rather look like
@@
The same relates to the bundle $Z$, which is now redefined as the product of two copies of the bundle of symmetric $(0,2)$-tensors over $S$. Similarly, we redefine in an obvious way all the decompositions and the projections introduced just above, as well as the function $\xi$. Now we can give a proof of Lemma 4.5.
Proof of Lemma 4.5..
Note that (4.5) and [34, Theorem II.2.3.1] imply that $\mathcal{E}_{+}$ and $\mathcal{E}$ are tame Fréchet manifolds. The same reference implies that so is $\mathcal{S}$.
By Lemma 4.2, there exists a neighborhood $U_{e}$ of $e$ in $\mathcal{E}$ such that for each $e^{\prime}\in U$, the differential $d_{e^{\prime}}\mathcal{I}$ is injective. We assume further that $U_{e}$ is sufficiently small so that the decomposition (4.6) holds and so that for every $e^{\prime}\in U_{e}$ the bundle morphisms $\alpha_{e^{\prime},1}$, $\alpha_{e^{\prime}_{2}}$, $\beta_{e^{\prime},1}$ and $\beta_{e^{\prime},2}$ are isomorphisms. Using (4.6) and decompositions $V=V_{1}\times V_{2}$ and $Z=Z_{1}\times Z_{2}$, we view the differential as a map
@@
By Lemma 4.7, each operator $P_{e^{\prime}}$ is elliptic of index $-(12k-12)$. By assumption, for each $e^{\prime}\in U_{e}$, the map $d_{e^{\prime}}\mathcal{I}$ is injective, hence, so is $\mathcal{D}_{1,e^{\prime}}:=\mathcal{D}_{1}(e^{\prime},.,.)$. It is well-known that ${\rm dim}\mathcal{R}_{F}=6k-6$, see, e.g. [22]. Because $H^{1}_{\rho}(\mathfrak{g})$ is the tangent space to $\mathcal{R}\cong\mathcal{R}_{F}\times\mathcal{R}_{F}$, we have ${\rm dim}H^{1}_{\rho}(\mathfrak{g})=12k-12$. It follows that for each $e^{\prime}\in U_{e}$, the map $\mathcal{D}_{1,e^{\prime}}$ is an isomorphism. Now by [34, Theorem II.3.3.3], the inverses to $\mathcal{D}_{1,e^{\prime}}$ constitute a smooth tame family of linear maps.
For each $e^{\prime}\in U_{e}$ and $\tau\in H^{1}_{\rho}(\mathfrak{g})$ the second part $\mathcal{D}_{2}(e^{\prime},\tau,.)$ is induced by a bundle diffeomorphism $V_{2}\rightarrow Z_{2}$, which depends smoothly on $e^{\prime}$ and $\tau$. Altogether this means that the inverses to $d\mathcal{I}$ constitute a smooth tame family of linear maps.
It follows that we can apply the Nash–Moser inverse function theorem [34, Theorem III.1.1.1], which implies that $\mathcal{I}$ is a local homeomorphism around $e$ onto an open set. ∎

### v2

@@ 4.2. Local infinitesimal rigidity
Similarly we define a differential operator $\mathcal{L}$. Tamburelli showed that for all such $Z^{\prime}$ the operators $\mathcal{L}_{i}$, $\mathcal{L}$, as operators from $C^{\infty}(Z^{\prime})$ to $C^{\infty}(\Lambda^{2}TS\otimes TS)$, are elliptic of index 6k66k-6 with trivial cokernel, where $k$ is the Euler characteristic of $S$. See the proof of Lemma 3.2 in [63]. (Compare also with [42, Lemma 3.1] of Labourie.) Hence every $\mathcal{L}_{i}$ on $C^{\infty}(Z_{i})$ has kernel of dimension 6k66k-6, where $C^{\infty}(Z_{i})$ is the tangent space to $C^{\infty}(J)$ at $a_{i}$. We denote this kernel by $X_{i}$. The same holds for the kernel $X$ of $\mathcal{L}$ at $C^{\infty}(Z)$, the tangent space at $a$.
We claim that it is enough to show that $X_{i}$ converge to $X$ as subspaces of $\mathcal{A}$. Indeed, GC-fields for $s_{i}$ are those $b\in C^{\infty}(G_{i})$ that satisfy $d^{\nabla_{i}}b=0$. Hence, $T_{b_{i}}\mathcal{GC}(s_{i})=\mathcal{K}_{i}(X_{i})$, as well as $T_{b}\mathcal{GC}(s)=\mathcal{K}(X)$. Since $\mathcal{K}_{i}$ are isomorphisms converging to $\mathcal{K}$, if $X_{i}$ converge to $X$, then $T_{b_{i}}\mathcal{GC}(s_{i})$ converge to $T_{b}\mathcal{GC}(s)$. To show the convergence of $X_{i}$, we construct a sequence of bundle automorphisms of $T^{1}_{1}S$ converging to the identity that send $Z_{i}$ to $Z$. To this purpose, we show that there exists a sequence of projectivized operator fields $p_{i}\in C^{\infty}(\mathbb{P}(T^{1}_{1}S))$, $C^{\infty}$-converging to the identity, such that $p_{i}a_{i}p_{i}^{-1}=a$. Here we note that the conjugation operation is well-defined for invertible elements of the projectivized operator bundle.
For the moment, let us consider a vector space $V\cong\mathbb{R}^{2}$ with some orientation. Let $J$ be the space of orientation-preserving almost-complex structures on $V$. Pick $a_{1}\in J, aa$$\neq a_{1}$. We claim that there exists a canonical $p\in{\rm PSL}(2,\mathbb{R})$ such that $pa_{1}p^{-1}=a$, where by canonical we mean that it is independent on a choice of basis or on a choice of any additional structure. Indeed, under the identification ${\rm PSL}(2,\mathbb{R})\cong\mathbb{A}d\mathbb{S}^{3}$ the space $J$ becomes the dual plane to the identity, or, in other words, the set of points at timelike distance $\pi/2$ from the identity. Under the identification
@@
the diagonal is the set of isometries fixing the identity. Hence, there exists a unique isometry fixing the identity, sending $a_{1}$ to $a$ and preserving the geodesic between them. This isometry is given by conjugation by $p\in{\rm PSL}(2,\mathbb{R})$.
Now we return to our setting. For every $i$ at every point of $S$ we pick the respective element of the fiber of $\mathbb{P}(T^{1}_{1}S)$ conjugating $a_{i}$ to $a$. This produces a smooth section $p_{i}$ of the projectivized bundle $\mathbb{P}(T^{1}_{1}S)$ such that $p_{i}a_{i}p_{i}^{-1}=a$.
Denote now $W^{1,2}(Z)$ by $\mathcal{Z}$. The conjugation by the projectivized operator fields $p_{i}$ define isomorphisms $\mathcal{R}_{i}:\mathcal{A}\rightarrow\mathcal{A}$ that converge to the identity in the sense of operator norms. The operators $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ are Fredholm operators from $\mathcal{Z}$ to $L^{2}(\Lambda^{2}TS\otimes TS)$ that converge to a Fredholm operator $\mathcal{L}$. All $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ and $\mathcal{L}$ have the same dimension of kernel and zero cokernel. Denote the kernels of $(\mathcal{R}_{i})_{*}\mathcal{L}_{i}$ by $Y_{i}$. For every $a^{\prime}\in\mathcal{A}$ we have
$d(a^{\prime},X)\gamma(\mathcal{L})\leq\|\mathcal{L}a^{\prime}\|.$
Since $\mathcal{L}$ is Fredholm, its range is closed, thus $\gamma(\mathcal{L})>0$. It follows that for $a^{\prime}\in Y_{i}$, $\|a^{\prime}\|\leq 1$, we get
$d(a^{\prime},X)\leq\gamma^{-1}(\mathcal{L})\|(\mathcal{L}-(\mathcal{R}_{i})_{*}\mathcal{L}_{i})\|.$
@@
We can now prove Proposition 4.2.
Proof of Proposition 4.2..
@@
We can assume that the sequences $\dot{b}_{i,+}$, $\dot{b}_{i,-}$ are uniformly bounded. From the mentioned results of Tamburelli, $T_{b_{+}}\mathcal{GC}_{+}(s_{+})$, $T_{b_{-}}\mathcal{GC}_{-}(s_{-})$ are finite-dimensional. Hence, by a combination of Lemma 4.4 and Lemma 4.1, these sequences subconverge to $\dot{b}_{+}\in T_{b_{+}}\mathcal{GC}_{+}(s_{+})$ and $\dot{b}_{-}\in T_{b_{-}}\mathcal{GC}_{-}(s_{-})$. Due to the definition of the maps $\phi_{s_{\pm},\pm}$ via formula (2.4), equation (4.2) implies that
@@ 4.3. Local homeomorphism
Now our goal is to prove the following statement:
Proposition 4.5.
Let $e\in\mathcal{E}_{F}$. There exists a neighborhood $U_{e}$ of $e$ in $\mathcal{E}$ such that the image of $U_{e}$ by $\mathcal{I}$ is open and $\mathcal{I}|_{U_{e}}$ is a homeomorphism onto its image.
We need a preliminary work. Let $\Sigma$ be a spacelike surface in $\mathbb{A}d\mathbb{S}^{3}$, $n$ be the field of future-oriented unit normals, and $u$ be a vector field over $\Sigma$ (not necessarily tangent), which is decomposed as $u=v+an$ for a tangent vector field $v$ and a function $a$. Denote the first and the second fundamental forms of $\Sigma$ by $\mathrm{I}$ and $\mathrm{I\!I}$ respectively, and denote by $\nabla$ the induced connection on $\Sigma$. It is a standard computation, to check that the derivative of $\mathrm{I}$ along $u$ is given by (see, e.g., [32])
@@
The bundle $Z$ of symmetric $(0,2)$-tensors on $\Sigma$ has rank 3. Denote by $Z_{1}$ its subbundle of traceless tensors with respect to $\mathrm{I\!I}$ and denote $V_{1}:=T\Sigma$. Note that the elements of $Z_{1}$ are orthogonal to $\mathrm{I\!I}$ with respect to $\mathrm{I\!I}$. Furthermore, let $b$ be the shape operator, $j$ be the almost-complex structure of $\mathrm{I\!I}$ and $\nabla^{\mathrm{I\!I\!I}}$ be the Levi–Civita connection of the third fundamental form. Define a differential operator $\mathcal{P}:C^{\infty}(V_{1})\rightarrow C^{\infty}(Z)$ by
@@
The significance of $\mathcal{P}$ is in that the kernel of $\mathcal{P}$ is exactly the space of the tangent components to the isometric vector fields on $\Sigma$. A direct computation (see, e.g., [57, Proposition 8.4]) shows that its image is in $C^{\infty}(Z_{1})$. Furthermore, it was proven in [57, Proposition 8.4] that
@@
Denote by $Z_{2}$ the subbundle of $Z$ spanned by $\mathrm{I\!I}$. Let $V_{2}$ be the normal subbundle of $T\mathbb{A}d\mathbb{S}^{3}$ over $\Sigma$. Formula (4.3) establishes a natural isomorphism
(4.5) $Q:V_{2}\rightarrow Z_{2}.$
When $f:\widetilde{S}\rightarrow\mathbb{A}d\mathbb{S}^{3}$ is an equivariant spacelike embedding, we can pullback $\mathcal{P}$ to $S$. It has the same principal symbol as a similarly defined operator, obtained by replacing $\nabla^{\mathrm{I\!I\!I}}$ with $\nabla^{\mathrm{I\!I}}$, the Levi–Civita connection of $\mathrm{I\!I}$. The latter operator has the same kernel and cokernel as the standard Cauchy–Riemann operator $\bar{\partial}$ (defined on de Rham complex tensored with $TS$). It follows from the Atiyah–Singer theorem (see, e.g., [34, Section 2]) that
@@
For a closed surface $S$ of genus $k$, as an operator $C^{\infty}(TS)\rightarrow C^{\infty}(Z_{1})$, where the latter is now considered as a bundle over $S$, $\mathcal{P}$ is an elliptic operator of index (6k6)-(6k-6).
[…]
(4.6) $T_{e}\mathcal{E}_{+}\cong H^{1}_{\rho}(\mathfrak{g})\times C^{\infty}(V)$
via
$T_{e}\mathcal{E}_{+}\cong C^{\infty}_{\rho,n}(\widetilde{V})={\rm Im}(\nu)\times{\rm Ker}(\zeta)\cong H^{1}_{\rho}(\mathfrak{g})\times C^{\infty}(V).$
Consider now a neighborhood $U$ of $e$ in $\mathcal{E}_{+}$ and a chart $\phi:U\rightarrow T_{e}\mathcal{E}_{+}$ adapted for the projection $\pi_{+}:\mathcal{E}_{+}\rightarrow\mathcal{R}$. Here “adapted” means that for any $e^{\prime}\in U$ and any $\tau\in H^{1}_{\rho}(\mathfrak{g})$ the composition
@@
is nonzero. It is possible to choose such a chart since $\pi_{+}$ is a submersion. We get an identification
@@
Every $e^{\prime}\in U$ defines a decomposition of $V$ into two subbundles, the tangent one, and the normal one (with respect to the metric on $V$ induced by $e^{\prime}$). We denote them by $V_{e^{\prime},1}$ and $V_{e^{\prime},2}$ respectively. Define $V_{1}:=V_{e,1}$ and $V_{2}:=V_{e,2}$. Let $\alpha_{e^{\prime},1}:V_{e^{\prime},1}\rightarrow V_{1}$ and $\alpha_{e^{\prime},2}:V_{e^{\prime},2}\rightarrow V_{2}$ be the orthogonal projections with respect to the metric on $V$ induced by $e$.
Also, every $e^{\prime}\in U$ defines a decomposition of $Z$ into two subbundles, $Z_{e^{\prime},1}$ and $Z_{e^{\prime},2}$, where $Z_{e^{\prime},2}$ is spanned by $\mathrm{I\!I}_{e^{\prime}}$, the second fundamental form induced on $S$ by $e^{\prime}$, and $Z_{e^{\prime},1}$ is $\mathrm{I\!I}_{e^{\prime}}$-orthogonal to it. Define $Z_{1}:=Z_{e,1}$ and $Z_{2}:=Z_{e,2}$. Let $\beta_{e^{\prime},1}:Z_{e^{\prime},1}\rightarrow Z_{1}$ and $\beta_{e^{\prime},2}:Z_{e^{\prime},2}\rightarrow Z_{2}$ be the orthogonal projections with respect to the metric on $Z$ induced by $\mathrm{I\!I}_{e}$. For the map $\xi:U\times H^{1}_{\rho}(\mathfrak{g})\rightarrow C^{\infty}(Z)$ defined above, denote its projections to $Z_{1}$ and to $Z_{2}$ by $\xi_{1}$ and $\xi_{2}$.
Now, however, we have to slightly change all this notation. Namely, we now pick $e\in\mathcal{E}$ rather than from $\mathcal{E}_{+}$. Then for some neighborhood $U$ of $e$ decomposition (4.7) will rather look like
@@
The same relates to the bundle $Z$, which is now redefined as the product of two copies of the bundle of symmetric $(0,2)$-tensors over $S$. Similarly, we redefine in an obvious way all the decompositions and the projections introduced just above, as well as the function $\xi$. Now we can give a proof of Proposition 4.5.
Proof of Proposition 4.5..
Note that (4.6) and [36, Theorem II.2.3.1] imply that $\mathcal{E}_{+}$ and $\mathcal{E}$ are tame Fréchet manifolds. The same reference implies that so is $\mathcal{S}$.
By Proposition 4.2, there exists a neighborhood $U_{e}$ of $e$ in $\mathcal{E}$ such that for each $e^{\prime}\in U_{e}$, the differential $d_{e^{\prime}}\mathcal{I}$ is injective. We assume further that $U_{e}$ is sufficiently small so that decomposition (4.7) holds and so that for every $e^{\prime}\in U_{e}$ the bundle morphisms $\alpha_{e^{\prime},1}$, $\alpha_{e^{\prime}_{2}}$, $\beta_{e^{\prime},1}$ and $\beta_{e^{\prime},2}$ are isomorphisms. Using (4.7) and decompositions $V=V_{1}\times V_{2}$ and $Z=Z_{1}\times Z_{2}$, we view the differential as a map
@@
From Lemma 4.6 and the definition of $\xi$, it follows that indeed $\mathcal{D}_{1}\times\mathcal{D}_{2}=d\mathcal{I}$.
By Lemma 4.7, each operator $P_{e^{\prime}}$ is elliptic of index (12k12)-(12k-12). By assumption, for each $e^{\prime}\in U_{e}$, the map $d_{e^{\prime}}\mathcal{I}$ is injective, hence, so is $\mathcal{D}_{1,e^{\prime}}:=\mathcal{D}_{1}(e^{\prime},\cdot,\cdot)$. It is well-known that ${\rm dim}\mathcal{R}_{F}=6k-6$, see, e.g. [24]. Because $H^{1}_{\rho}(\mathfrak{g})$ is the tangent space to $\mathcal{R}\cong\mathcal{R}_{F}\times\mathcal{R}_{F}$, we have ${\rm dim}H^{1}_{\rho}(\mathfrak{g})=12k-12$. Hence, the index of $\mathcal{D}_{1,e^{\prime}}$ is zero. It follows that for each $e^{\prime}\in U_{e}$, the map $\mathcal{D}_{1,e^{\prime}}$ is an isomorphism. Now by [36, Theorem II.3.3.3], the inverses to $\mathcal{D}_{1,e^{\prime}}$ constitute a smooth tame family of linear maps.
For each $e^{\prime}\in U_{e}$ and $\tau\in H^{1}_{\rho}(\mathfrak{g})$, $\mathcal{D}_{2}(e^{\prime},\tau,\cdot)$ is induced by a bundle diffeomorphism $V_{2}\rightarrow Z_{2}$, which depends smoothly on $e^{\prime}$ and $\tau$. Altogether this means that the inverses to $d\mathcal{I}$ constitute a smooth tame family of linear maps.
It follows that we can apply the Nash–Moser inverse function theorem [36, Theorem III.1.1.1], which implies that $\mathcal{I}$ is a local homeomorphism around $e$ onto an open set. ∎